### v1

@@ 6 Connecting the counting hierarchy to the $\tau$-conjecture
Instead of taking the Pochhammer polynomial, there are many other possible choices for some explicit polynomial, see [9].
@@ 7 Preliminaries II: The VW-hierarchy
In this section, we define different variants of the $\mathsf{VW}$-hierarchy, which will be analogous to #W\#W-hierarchy, see [4]. We will consider circuits that can have unbounded fanin gates.
@@ 8 Conditional collapsing of VW-hierarchy and applications
Let us recall the definition of $k$-degree $n$-variate $(n\geq k)$ elementary symmetric polynomial $\sigma_{n,k}(\mathbf{X})\;:=\;\sum_{y\in\langle{n\atop{k}}\rangle}\,X_{1}^{y_{1}}X_{2}^{y_{2}}\dots X_{n}^{y_{n}}\;.$ It is known that $(\sigma_{n,k})_{n}\in\mathsf{VP}^{0}$, with a simple dynamic programming algorithm; see [17, Section 4]. Let us define a new polynomial family $B_{n,k}(\mathbf{X})$, which will be important in the latter part of the section: $B_{n,k}(\mathbf{X})\;:=\;\sum_{t=0}^{n-k}\,(-1)^{t}\binom{k+t}{k}\cdot\sigma_{n,k+t}(\mathbf{X})\;$. The following claim is crucial:
@@
The above theorem holds in the unbounded regime as well, i.e., $\mathsf{VW}_{\textup{nb}}^{0}[\mathsf{P}]\;\neq\;\mathsf{VFPT}_{\textup{nb}}^{0}\;\implies\;\mathsf{VP}_{\textup{nb}}^{0}\;\neq\;\mathsf{VNP}_{\textup{nb}}^{0}$ (which further implies $\mathsf{VP}^{0}\neq\mathsf{VNP}^{0}$, see [20]). Similarly, $\mathsf{VW}_{\textup{nb}}[\mathsf{P}]\;\neq\;\mathsf{VFPT}_{\textup{nb}}\;\implies\;\mathsf{VP}_{\textup{nb}}\;\neq\;\mathsf{VNP}_{\textup{nb}}$.
@@ 10 Hardness of the restricted permanent
Our goal is to write $P$ as an fpt-projection of a $(k,c)$-restricted permanent on a $(c,b)$-nice graph $H$ for certain constants $c$ and $b$. The construction will have two main components. One corresponds to the formula $F$, the other one is similar to the rosetta graph in Valiant’s proof of the $\mathsf{\#P}$-hardness of the permanent, see e.g. [8]. The first component will be the bounded treewidth part of $H$, the second one will be the high girth part.
@@ 10.1 The graph $G_{1}$
We first design a graph $G_{1}$ from $F$. $G_{1}$ will have iff-coupled edges (pairs of edges). When we have a pair of iff-coupled edges, then either both of them appear in a cycle cover or none of them, see also [8]. Later, we will enforce this by connecting iff-coupled edges with appropriate gadgets.
@@
The graph will be constructed in an iterative manner, adding new nodes and edges gradually. Each parse tree of $F$ will correspond to one consistent cycle cover and vice versa. Recall that a parse tree of an algebraic formula is a subtree that contains the root, for each multiplication gate it contains all children and for each addition gate it contains exactly one child, see [22].
@@ 10.4 The graph $G_{3}$
Now the graph $G_{3}$ is obtained by replacing the iff-couplings with the gadget in Figure 6. If the two edge $(x,y)$ and $(u,v)$ are iff-coupled, then we subdivide the edges with the nodes $a$ and $b$ and connect them with the subgraph as depicted. For each iff-coupling, we insert a new subgraph. If we do not write a weight explicitly, then the weight of the edge is $1$. Similar gadgets were developed in the past, see e.g. [11]. The difference in our gadget is that we have a $4$-cycle between $a$ and $b$ instead of a $2$-cycle. This will be crucial, since $(k,c)$-restricted cycle covers are sensitive to changes of cycle lengths.
@@
If the cycle cover $C$ is $(2k,2)$-restricted, the the resulting cycle covers will be $(3k,4)$-restricted, since each iff-coupled edge is subdivided. Each edge is only subdivided once except for the loops at the input gates that were labeled with a $Y$-variable. These are subdivided twice, yielding a cycle of length $3$. Furthermore, the internal cycles of the iff-gadgets have length at most $4$.
@@ 11 Upper bound
Bodlaender and Hagerup [7] show that whenever a graph has bounded tree-width, then there is a binary tree-decomposition of logarithmic height. Moreover, we can assume that the tree decomposition is nice, see e.g. [10] for a definition, and the height is still logarithmic.

### v2

@@ 6 Connecting the counting hierarchy to the $\tau$-conjecture
Instead of taking the Pochhammer polynomial, there are many other possible choices for some explicit polynomial, see [BÜR09].
@@ 7 Preliminaries II: The VW-hierarchy
In this section, we define different variants of the $\mathsf{VW}$-hierarchy, which will be analogous to #W\#W-hierarchy, see [BE19]. We will consider circuits that can have unbounded fanin gates.
@@ 8 Conditional collapsing of VW-hierarchy and applications
Let us recall the definition of $k$-degree $n$-variate $(n\geq k)$ elementary symmetric polynomial $\sigma_{n,k}(\mathbf{X})\;:=\;\sum_{y\in\langle{n\atop{k}}\rangle}\,X_{1}^{y_{1}}X_{2}^{y_{2}}\dots X_{n}^{y_{n}}\;.$ It is known that $(\sigma_{n,k})_{n}\in\mathsf{VP}^{0}$, with a simple dynamic programming algorithm; see [JS16, Section 4]. Let us define a new polynomial family $B_{n,k}(\mathbf{X})$, which will be important in the latter part of the section: $B_{n,k}(\mathbf{X})\;:=\;\sum_{t=0}^{n-k}\,(-1)^{t}\binom{k+t}{k}\cdot\sigma_{n,k+t}(\mathbf{X})\;$. The following claim is crucial:
@@
The above theorem holds in the unbounded regime as well, i.e., $\mathsf{VW}_{\textup{nb}}^{0}[\mathsf{P}]\;\neq\;\mathsf{VFPT}_{\textup{nb}}^{0}\;\implies\;\mathsf{VP}_{\textup{nb}}^{0}\;\neq\;\mathsf{VNP}_{\textup{nb}}^{0}$ (which further implies $\mathsf{VP}^{0}\neq\mathsf{VNP}^{0}$, see [KP11]). Similarly, $\mathsf{VW}_{\textup{nb}}[\mathsf{P}]\;\neq\;\mathsf{VFPT}_{\textup{nb}}\;\implies\;\mathsf{VP}_{\textup{nb}}\;\neq\;\mathsf{VNP}_{\textup{nb}}$.
@@ 10 Hardness of the restricted permanent
Our goal is to write $P$ as an fpt-projection of a $(k,c)$-restricted permanent on a $(c,b)$-nice graph $H$ for certain constants $c$ and $b$. The construction will have two main components. One corresponds to the formula $F$, the other one is similar to the rosetta graph in Valiant’s proof of the $\mathsf{\#P}$-hardness of the permanent, see e.g. [BÜR00]. The first component will be the bounded treewidth part of $H$, the second one will be the high girth part.
@@ 10.1 The graph $G_{1}$
We first design a graph $G_{1}$ from $F$. $G_{1}$ will have iff-coupled edges (pairs of edges). When we have a pair of iff-coupled edges, then either both of them appear in a cycle cover or none of them, see also [BÜR00]. Later, we will enforce this by connecting iff-coupled edges with appropriate gadgets.
@@
The graph will be constructed in an iterative manner, adding new nodes and edges gradually. Each parse tree of $F$ will correspond to one consistent cycle cover and vice versa. Recall that a parse tree of an algebraic formula is a subtree that contains the root, for each multiplication gate it contains all children and for each addition gate it contains exactly one child, see [MP08].
@@ 10.4 The graph $G_{3}$
Now the graph $G_{3}$ is obtained by replacing the iff-couplings with the gadget in Figure 6. If the two edge $(x,y)$ and $(u,v)$ are iff-coupled, then we subdivide the edges with the nodes $a$ and $b$ and connect them with the subgraph as depicted. For each iff-coupling, we insert a new subgraph. If we do not write a weight explicitly, then the weight of the edge is $1$. Similar gadgets were developed in the past, see e.g. [DHM+14]. The difference in our gadget is that we have a $4$-cycle between $a$ and $b$ instead of a $2$-cycle. This will be crucial, since $(k,c)$-restricted cycle covers are sensitive to changes of cycle lengths.
@@
If the cycle cover $C$ is $(2k,2)$-restricted, the resulting cycle covers will be $(3k,4)$-restricted, since each iff-coupled edge is subdivided. Each edge is only subdivided once except for the loops at the input gates that were labeled with a $Y$-variable. These are subdivided twice, yielding a cycle of length $3$. Furthermore, the internal cycles of the iff-gadgets have length at most $4$.
@@ 11 Upper bound
Bodlaender and Hagerup [BH95] show that whenever a graph has bounded tree-width, then there is a binary tree-decomposition of logarithmic height. Moreover, we can assume that the tree decomposition is nice, see e.g. [CFK+15] for a definition, and the height is still logarithmic.